OpenStudy (anonymous):

$2895 is invested into a savings account and compounded continuously. What interest rate is required for the investment to triple in 12 years? PLEASE HELP

OpenStudy (anonymous):

A=Pe^rt formula

OpenStudy (danjs):

using the formula you gave, and assuming the rate compounds once per year, P=2895 A=3*P = 8685 t=3 compounded per year or t=36 (compounded monthly) solve for the rate r.